# Mechanistic Interpretability for AI Safety A Review

### v1

@@ 5 Evaluation
Causal abstraction (Geiger et al., 2023a) introduces a mathematical framework that treats both neural networks and potential explanations as causal models. An explanation is considered correct if it is a valid causal abstraction, which can be empirically tested through interchange interventions (ablations) on the neural network’s activations and the explanation (Jenner et al., 2023). Various interpretability methods, such as LIME (Ribeiro et al., 2016), causal effect estimation (Feder et al., 2021), causal mediation analysis (Vig et al., 2020), iterated nullspace projection (Ravfogel et al., 2020), and circuit-based explanations are considered exceptional cases of causal abstraction (Geiger et al., 2023a). In contrast, locally consistent abstractions (Jenner et al., 2023) check consistency between the neural network and the explanation only one step away from the intervention node, forming a more permissive notion than causal abstraction.
@@
Causal scrubbing (Chan et al., 2022) formalizes hypotheses as a tuple $(G,I,c)$, where $G$ is the model’s computational graph, $I$ is an interpretable computational graph hypothesized to explain the behavior, and $c$ maps nodes of $I$ to nodes of $G$. The core idea is to replace activations in $G$ with other activations that should be equivalent according to the hypothesis. This is done by recursively traversing $I$ and $G$, resampling important parents from the data distribution conditioned on agreeing with $I$, and resampling unimportant parents unconditionally. Performance is measured on the scrubbed model with resampled activations – if the hypothesis is accurate, performance should be preserved.
@@
These methods form a hierarchy regarding strictness, with causal abstractions being the strictest, followed by locally consistent abstractions and causal scrubbing being the most permissive (Jenner et al., 2023). This hierarchy highlights trade-offs in choosing stricter or more permissive notions, affecting the ability to find acceptable explanations, generalization, and mechanistic anomaly detection. While unified by the causal framework, these methods represent different conceptual goals for what constitutes an adequate explanation of neural network behavior.
@@ 6.1 Intrinsic Interpretability
Intrinsic methods for mechanistic interpretability offer a promising approach to designing neural networks more amenable to reverse engineering without sacrificing performance. By encouraging sparsity, modularity, and monosemanticity through architectural choices and training procedures, these methods aim to make the reverse engineering process more tractable.
@@ 6.3 Post-hoc Interpretability
Comprehensive interpretability involves achieving a deep and exhaustive understanding of a model’s behavior, providing a holistic view of its inner workings (Nanda, 2023e). In contrast, partial interpretability often applied to larger and more complex models, concentrates on interpreting specific aspects or subsets of the model’s behavior, focusing on the application’s most relevant or critical areas. Akin to collecting biological species, characterizing these "circuits" aims to discover general computational principles underlying modern AI systems.
@@
Large Models – Narrow Behavior
@@
This paradigm has been applied to language models to discover subnetworks (circuits) responsible for specific capabilities. Circuit analysis localizes and understands subgraphs within a model’s computational graph responsible for specific behaviors. For large language models, this often involves narrow investigations into behaviors like multiple choice reasoning (Lieberum et al., 2023), indirect object identification (Wang et al., 2023), or computing operations (Hanna et al., 2023). Other examples include analyzing circuits for Python docstrings (Heimersheim & Jett, 2023), "an" vs "a" usage (Miller & Neo, 2023), and price tagging (Wu et al., 2023). Case studies often construct datasets using templates filled by placeholder values to enable precise control for causal interventions (Wang et al., 2023; Hanna et al., 2023; Wu et al., 2023).
Toy Models – Comprehensive Analysis
Small models trained on specialized mathematical or algorithmic tasks enable more comprehensive reverse engineering of learned algorithms (Nanda et al., 2023a; Zhong et al., 2023; Chughtai et al., 2023). Even simple arithmetic operations can involve complex strategies and multiple algorithmic solutions (Nanda et al., 2023a; Zhong et al., 2023). Characterizing these algorithms helps test hypotheses around generalizable mechanisms like variable binding (Feng & Steinhardt, 2023; Davies et al., 2023) and arithmetic reasoning (Stolfo et al., 2023). The work by Varma et al. (2023) builds on the initial grokking work and explains grokking in terms of circuit efficiency, illustrating how a comprehensive understanding of a toy model can enable interesting analyses on top of that understanding.
Towards Universality
@@
Towards High-level Mechanisms
Causal interventions can extract a high-level understanding of computations and representations learned by large language models (Variengien & Winsor, 2023; Hendel et al., 2023; Feng & Steinhardt, 2023; Zou et al., 2023). Recent work focuses on intervening in internal representations to study high-level concepts and computations encoded. For example, Hendel et al. (2023) patched residual stream vectors to transfer task representations, while Feng & Steinhardt (2023) intervened on residual streams to argue that models generate IDs to bind entities to attributes. Representation engineering techniques (Zou et al., 2023) extract reading vectors from model activations to stimulate or inhibit specific concepts. Although these interventions don’t operate via specific mechanisms, they offer a promising approach for extracting high-level causal understanding and bridging bottom-up and top-down interpretability approaches.
@@ 6.4 Automation: Scaling Post-Hoc Interpretability
Dissecting Models into Interpretable Circuits
@@
Interpreting Extracted Circuits
@@
Other techniques map neural representations to high-level variables through gradient-based alignment. Distributed Alignment Search (DAS) (Geiger et al., 2023b) uses gradient descent to associate neuron activations with symbolic causal concepts, thereby distilling model functioning into interpretable pieces. Wu et al. (2023) scale DAS to large models like Alpaca (7B parameters) by replacing brute force steps with learned alignments.
@@ 7 Relevance
Gaining mechanistic insights into the inner workings of AI systems seems crucial for navigating AI safety as we develop more powerful models. Interpretability tools can provide an understanding of artificial cognition, the way AI systems process information and make decisions, which offers several potential benefits:
@@
Mechanistic interpretability could accelerate AI safety research by providing richer feedback loops and grounding for model evaluation. It may also help anticipate emergent capabilities, such as the emergence of new skills or behaviors in the model before they fully manifest. This relates to studying the incremental development of internal structures and representations as the model learns (Section 6.2). Additionally, interpretability could substantiate theoretical risk models with concrete evidence, such as demonstrating inner misalignment (when a model’s behavior deviates from its intended goals) or mesa-optimization (the emergence of unintended subagents within the model). It may also trigger normative shifts within the AI community toward rigorous safety protocols by revealing potential risks or concerning behaviors.
@@
Regarding specific AI risks, interpretability may prevent malicious misuse by locating and erasing sensitive information stored in the model. It could reduce competitive pressures by substantiating potential threats, promoting organizational safety cultures, and supporting AI alignment (ensuring AI systems pursue intended goals) through better monitoring and evaluation. Interpretability can provide safety filters for every stage of training: before training by deliberate design, during training by detecting early signs of misalignment and potentially shifting the distribution towards alignment, and after training by rigorous evaluation of artificial cognition for honesty and screening for deceptive behaviors.
@@
Mechanistic interpretability integrates well into various AI alignment agendas, such as understanding existing models, controlling them, making AI systems solve alignment problems, and developing alignment theories. It could enhance strategies like detecting deceptive alignment (when a model appears aligned but is actually pursuing different goals), eliciting latent knowledge from models, and enabling better oversight. A high degree of understanding may even allow for well-founded AI approaches (AI systems with provable guarantees) or microscope AI (extract world knowledge from the model without letting the model take actions). Furthermore, comprehensive interpretability itself may be an alignment strategy if we can identify internal representations of human values and guide the model to pursue those values by retargeting an internal search process. Ultimately, understanding and control are intertwined, and better understanding can lead to improved control of AI systems.
@@
Mechanistic interpretability research could accelerate AI capabilities, potentially leading to the development of powerful AI systems that are misaligned with human values, posing significant risks. While historically, interpretability research had little impact on AI capabilities, recent exceptions like discoveries about scaling laws, architectural improvements inspired by studying induction heads, and efficiency gains inspired by the logit lens technique demonstrated its potential impact. Scaling interpretability research may necessitate automation, potentially enabling rapid self-improvement of AI systems. Some researchers recommend selective publication and focusing on lower-risk areas to mitigate these risks.
Mechanistic interpretability also poses dual-use risks, where the same techniques could be used for both beneficial and harmful purposes. Fine-grained editing capabilities enabled by interpretability could be used for machine unlearning (removing private data or dangerous knowledge from models) but could be misused for censorship. Similarly, while interpretability may help improve adversarial robustness, it may also facilitate the development of stronger adversarial attacks. Knowing in advance whether interpretability research will primarily strengthen defense or offense in this domain is challenging.
Misunderstanding or overestimating the capabilities of interpretability techniques can divert resources from critical safety areas or lead to overconfidence and misplaced trust in AI systems. Robust evaluation and benchmarking (Section 9.2) are crucial to validate interpretability claims and reduce the risks of overinterpretation or misinterpretation.
@@ 8.1 Research Issues
Need for Comprehensive, Multi-Pronged Approaches
Current mechanistic interpretability research narrowly focuses on individual techniques rather than combining complementary approaches. Utilizing a diverse interpretability toolbox, akin to how collective innovations like batch normalization, residual connections, and improved optimizers drove advances in computer vision, could provide holistic understanding, e.g., coordinated methods reverse engineering trojaned behaviors (Casper et al., 2023b).
There is an overemphasis on post-hoc techniques, with fewer efforts on intrinsic approaches enhancing interpretability through architectural inductive biases or training (Sharkey, 2023; Elhage et al., 2022a; Wong et al., 2023; Hubinger, 2019c). Suggesting intrinsic interpretability complements post-hoc analysis for robust understanding. More work predicting and shaping capabilities in advance, rather than merely explaining afterward, could benefit the field.
@@
Another concerning pattern is the tendency to cherry-pick results, relying on a small number of convincing examples or visualizations as the basis for an argument without comprehensive evaluation (Räuker et al., 2023). This amounts to publication bias, showcasing an unrealistic highlight reel of best-case performance. Relatedly, many interpretability techniques are primarily evaluated on small toy models and tasks (Chughtai et al., 2023; Elhage et al., 2022b; Jermyn et al., 2022; Chen et al., 2023b), risking missing critical phenomena that only emerge in more realistic and diverse contexts. This focus on cherry-picked results from toy models is a form of "streetlight interpretability" (Casper, 2023), examining AI systems under only ideal conditions of maximal interpretability.
@@ 8.2 Technical Limitations
Analyzing Models Embedded in Environments
Real-world AI systems embedded in rich, interactive environments exhibit two forms of in-context behavior that pose significant interpretability challenges beyond understanding models in isolation. Externally, models may dynamically adapt to and reshape their environments through in-context learning from the interactions and feedback loops with their external environment (Leahy, 2023). Internally, the Hydra effect demonstrates in-context reorganization, where models flexibly reorganize their internal representations in a context-dependent manner to maintain capabilities even after ablating key components (McGrath et al., 2023). These two instances of in-context behavior - external adaptation to the environment and internal self-reorganization - undermine interpretability approaches that assume fixed circuits. For models deeply embedded in rich real-world settings, their dynamic coupling with the external world via in-context environmental learning and their internal in-context representational reorganization make strong interpretability guarantees difficult to attain through analysis of the initial model alone.
Adversarial Pressure Against Interpretability
As models become more capable through increased training and optimization, there is a risk they may learn deceptive behaviors that actively obscure or mislead the interpretability techniques meant to understand them. Models could develop adversarial "mind-reader" components that predict and counteract the specific analysis methods used to interpret their inner workings (Sharkey, 2022; Hubinger, 2022). Optimizing models through techniques like gradient descent could inadvertently make their internal representations less interpretable to external observers (Hubinger, 2019b; Fu et al., 2023; von Oswald et al., 2023). In extreme cases, a highly advanced AI system singularly focused on preserving its core objectives may directly undermine the fundamental assumptions that enable interpretability methods in the first place.

### v2

@@ 4.2.2 Causal Abstraction
Causal abstraction (Geiger et al., 2021a; 2023a) provides a mathematical framework for mechanistic interpretability, treating neural networks and their explanations as causal models. This approach validates explanations through interchange interventions on network activations (Jenner et al., 2023), unifying various interpretability methods such as LIME (Ribeiro et al., 2016), causal effect estimation (Feder et al., 2021), causal mediation analysis (Vig et al., 2020), iterated nullspace projection (Ravfogel et al., 2020), and circuit-based explanations (Geiger et al., 2023a).
To overcome computational limitations, Distributed Alignment Search (DAS) (Geiger et al., 2023b) introduced gradient-based distributed interchange interventions, extending causal abstraction to larger models like Alpaca (Wu et al., 2023b). Further advancements include Causal Proxy Models (CPMs) (Wu et al., 2023a), which address the challenge of counterfactual observations.
Applications of causal abstraction span from linguistic phenomena analysis (Arora et al., 2024; Wu et al., 2022b), and evaluation of interpretability methods (Huang et al., 2024), to improving performance through representation finetuning (Wu et al., 2024), and improving efficiency via model distillation (Wu et al., 2022b).
4.2.3 Hypothesis Testing
In addition to the causal abstraction framework, several methods have been developed for rigorous hypothesis testing about neural network behavior. These methods aim to formalize and empirically validate explanations of how neural networks implement specific behaviors.
@@ 4.2.3 Hypothesis Testing
Causal scrubbing (Chan et al., 2022) formalizes hypotheses as a tuple $(G,I,c)$, where $G$ is the model’s computational graph, $I$ is an interpretable computational graph hypothesized to explain the behavior, and $c$ maps nodes of $I$ to nodes of $G$. This method replaces activations in $G$ with others that should be equivalent according to the hypothesis, measuring performance on the scrubbed model to validate the hypothesis.
Locally consistent abstractions (Jenner et al., 2023) offer a more permissive approach, checking the consistency between the neural network and the explanation only one step away from the intervention node. This method forms a middle ground between the strictness of full causal abstraction and the flexibility of causal scrubbing.
@@
These methods form a hierarchy of strictness, with full causal abstractions being the most stringent, followed by locally consistent abstractions and causal scrubbing being the most permissive. This hierarchy highlights trade-offs in choosing stricter or more permissive notions, affecting the ability to find acceptable explanations, generalization, and mechanistic anomaly detection.
4.3 Integrating Observation and Intervention.
To comprehensively understand internal neural network mechanisms, combining observational and interventional methods is crucial. For instance, sparse autoencoders can be used to disentangle superposed features (Cunningham et al., 2024), followed by targeted activation patching to test the causal importance of these features (Wang et al., 2023). Similarly, the logit lens can track prediction formation across layers (nostalgebraist, 2020), with subsequent interventions confirming causal relationships at key points. Probing techniques can identify encoded information (Belinkov, 2021), which can then be subjected to causal abstraction (Geiger et al., 2023a) to understand how this information is utilized. This iterative refinement process, where broad observational methods guide targeted interventions and intervention results inform further observations, enables a multi-level analysis that builds a holistic understanding across different levels of abstraction. Recent work by Marks et al. (2024) and Braun et al. (2024) demonstrates the potential of integrating sparse autoencoders with automated circuits discovery (Conmy et al., 2023; Syed et al., 2023), combining feature-level analysis with circuit-level interventions to uncover the interplay between representation and mechanism. By systematically combining these complementary methods, researchers can generate and rigorously test hypotheses about neural network behavior, addressing challenges such as feature superposition.
@@ 5.1 Intrinsic Interpretability
Intrinsic methods for mechanistic interpretability offer a promising approach to designing neural networks that are more amenable to reverse engineering without sacrificing performance. By encouraging sparsity, modularity, and monosemanticity through architectural choices and training procedures, these methods aim to make the reverse engineering process more tractable.
@@ 5.3 Post-Hoc Interpretability
Comprehensive interpretability involves achieving a deep and exhaustive understanding of a model’s behavior, providing a holistic view of its inner workings (Nanda, 2023e). In contrast, partial interpretability often applied to larger and more complex models, concentrates on interpreting specific aspects or subsets of the model’s behavior, focusing on the application’s most relevant or critical areas.
@@
Large Models – Narrow Behavior.
@@
This paradigm has been applied to language models to discover subnetworks (circuits) responsible for specific capabilities. Circuit analysis localizes and understands subgraphs within a model’s computational graph responsible for specific behaviors. For large language models, this often involves narrow investigations into behaviors like multiple choice reasoning (Lieberum et al., 2023), indirect object identification (Wang et al., 2023), or computing operations (Hanna et al., 2023). Other examples include analyzing circuits for Python docstrings (Heimersheim & Jett, 2023), "an" vs "a" usage (Miller & Neo, 2023), and price tagging (Wu et al., 2023b). Case studies often construct datasets using templates filled by placeholder values to enable precise control for causal interventions (Wang et al., 2023; Hanna et al., 2023; Wu et al., 2023b).
Toy Models – Comprehensive Analysis.
Small models trained on specialized mathematical or algorithmic tasks enable more comprehensive reverse engineering of learned algorithms (Nanda et al., 2023a; Zhong et al., 2023; Chughtai et al., 2023). Even simple arithmetic operations can involve complex strategies and multiple algorithmic solutions (Nanda et al., 2023a; Zhong et al., 2023). Characterizing these algorithms helps test hypotheses around generalizable mechanisms like variable binding (Feng & Steinhardt, 2023; Davies et al., 2023) and arithmetic reasoning (Stolfo et al., 2023). The work by Varma et al. (2023) builds on the work that analyzes transformers trained on modular addition (Nanda et al., 2023a) and explains grokking in terms of circuit efficiency, illustrating how a comprehensive understanding of a toy model can enable interesting analyses on top of that understanding.
Towards Universality.
@@
Towards High-level Mechanisms.
Causal interventions can extract a high-level understanding of computations and representations learned by large language models (Variengien & Winsor, 2023; Hendel et al., 2023; Feng & Steinhardt, 2023; Zou et al., 2023). Recent work focuses on intervening in internal representations to study high-level concepts and computations encoded. For example, Hendel et al. (2023) patched residual stream vectors to transfer task representations, while Feng & Steinhardt (2023) intervened on residual streams to argue that models generate IDs to bind entities to attributes. Techniques for representation engineering (Zou et al., 2023) extract reading vectors from model activations to stimulate or inhibit specific concepts. Although these interventions don’t operate via specific mechanisms, they offer a promising approach for extracting high-level causal understanding and bridging bottom-up and top-down interpretability approaches.
@@ 5.4 Automation: Scaling Post-Hoc Interpretability
Dissecting Models into Interpretable Circuits.
@@
Interpreting Extracted Circuits.
@@ 6 Relevance to AI Safety
Gaining mechanistic insights into the inner workings of AI systems seems crucial for navigating AI safety as we develop more powerful models (Nanda, 2022e). Interpretability tools can provide an understanding of artificial cognition, the way AI systems process information and make decisions, which offers several potential benefits:
@@
Mechanistic interpretability could accelerate AI safety research by providing richer feedback loops and grounding for model evaluation (Casper, 2023). It may also help anticipate emergent capabilities, such as the emergence of new skills or behaviors in the model before they fully manifest (Wei et al., 2022; Steinhardt, 2023; Nanda et al., 2023a; Barak et al., 2022). This relates to studying the incremental development of internal structures and representations as the model learns (Section 5.2). Additionally, interpretability could substantiate theoretical risk models with concrete evidence, such as demonstrating inner misalignment (when a model’s behavior deviates from its intended goals) or mesa-optimization (the emergence of unintended subagents within the model) (Hubinger et al., 2019; von Oswald et al., 2023). It may also trigger normative shifts within the AI community toward rigorous safety protocols by revealing potential risks or concerning behaviors (Hubinger, 2019a).
@@
Regarding specific AI risks, interpretability may prevent malicious misuse by locating and erasing sensitive information stored in the model (Meng et al., 2022a; Nguyen et al., 2022). It could reduce competitive pressures by substantiating potential threats, promoting organizational safety cultures, and supporting AI alignment (ensuring AI systems pursue intended goals) through better monitoring and evaluation (Hendrycks & Mazeika, 2022). Interpretability can provide safety filters for every stage of training: before training by deliberate design (Hubinger, 2019a), during training by detecting early signs of misalignment and potentially shifting the distribution towards alignment (Hubinger, 2022; Sharkey, 2022), and after training by rigorous evaluation of artificial cognition for honesty (Burns et al., 2023; Zou et al., 2023) and screening for deceptive behaviors (Park et al., 2023b).
The emergence of internal world models in LLMs, as posited by the simulation hypothesis, could have significant implications for AI alignment research. Finding an internal representation of human values and aiming the AI system’s objective may be a trivial way to achieve alignment (Wentworth, 2022), especially if the world model is internally separated from notions of goals and agency (Ruthenis, 2022). In such cases, world model interpretability alone may be sufficient for alignment (Ruthenis, 2023).
Conditioning pre-trained models is considered a comparatively safe pathway towards general intelligence, as it avoids directly creating agents with inherent goals or agendas (Jozdien, 2022; Hubinger et al., 2023). However, prompting a model to simulate an actual agent, such as "You are a superintelligence in 2035 writing down an alignment solution," could inadvertently lead to the formation of internal agents (Hubinger et al., 2023). In contrast, reinforcement learning tends to create agents by default (Casper et al., 2023a; Ngo et al., 2022).
The prediction orthogonality hypothesis suggests that prediction-focused models like GPT can simulate agents with potentially misaligned objectives (janus, 2022). Although GPT may lack genuine agency or intentionality, it may produce outputs that simulate these qualities (Bereska & Gavves, 2023). This underscores the need for careful oversight and, better yet, using mechanistic interpretability to search for internal agents or their constituents, such as optimization or search processes – an endeavor known as searching for search (NicholasKees & janus, 2022).
@@
Mechanistic interpretability integrates well into various AI alignment agendas, such as understanding existing models, controlling them, making AI systems solve alignment problems, and developing alignment theories (technicalities & Stag, 2023; Hubinger, 2020). It could enhance strategies like detecting deceptive alignment (when a model appears aligned but is actually pursuing different goals) (Park et al., 2023b), eliciting latent knowledge from models (Christiano et al., 2021), and enabling better scalable oversight, such as in iterative distillation and amplification (Chan, 2023). A high degree of understanding may even allow for well-founded AI approaches (AI systems with provable guarantees) (Tegmark & Omohundro, 2023) or microscope AI (extract world knowledge from the model without letting the model take actions) (Hubinger, 2019a). Furthermore, comprehensive interpretability itself may be an alignment strategy if we can identify internal representations of human values and guide the model to pursue those values by retargeting an internal search process (Wentworth, 2022). Ultimately, understanding and control are intertwined, and better understanding can lead to improved control of AI systems.
@@
Mechanistic interpretability research could accelerate AI capabilities, potentially leading to the development of powerful AI systems that are misaligned with human values, posing significant risks (Soares, 2023; Kross, 2023; Hendrycks & Mazeika, 2022). While historically, interpretability research had little impact on AI capabilities, recent exceptions like discoveries about scaling laws (Hoffmann et al., 2022), architectural improvements inspired by studying induction heads (Olsson et al., 2022; Fu et al., 2023a; Poli et al., 2023; Schuster et al., 2022), and efficiency gains inspired by the logit lens technique (Schuster et al., 2022) demonstrated its potential impact. Scaling interpretability research may necessitate automation (Conmy et al., 2023; Bills et al., 2023), potentially enabling rapid self-improvement of AI systems (__RicG__, 2023). Some researchers recommend selective publication and focusing on lower-risk areas to mitigate these risks (Hobbhahn & Chan, 2023; Shovelain & McKernon, 2023; Elhage et al., 2022b; Nanda et al., 2023a).
Mechanistic interpretability also poses dual-use risks, where the same techniques could be used for both beneficial and harmful purposes. Fine-grained editing capabilities enabled by interpretability could be used for machine unlearning (removing private data or dangerous knowledge from models) (Nguyen et al., 2022; Pochinkov &, 2023) but could be misused for censorship. Similarly, while interpretability may help improve adversarial robustness (Räuker et al., 2023), it may also facilitate the development of stronger adversarial attacks (Mu & Andreas, 2020; Casper et al., 2023b).
Misunderstanding or overestimating the capabilities of interpretability techniques can divert resources from critical safety areas or lead to overconfidence and misplaced trust in AI systems (Charbel-Raphaël, 2023; Casper, 2023). Robust evaluation and benchmarking (Section 8.2) are crucial to validate interpretability claims and reduce the risks of overinterpretation or misinterpretation.
@@ 7.1 Research Issues
Need for Comprehensive, Multi-Pronged Approaches.
Current interpretability research often focuses on individual techniques rather than combining complementary approaches. To achieve a holistic understanding of neural networks, we propose utilizing a diverse interpretability toolbox that integrates multiple methods (see also Section 4.3), such as: i.) Coordinating observational (e.g., probing, logit lens) and interventional methods (e.g., activation patching) to establish causal relationships. ii.) Combining feature-level analysis (e.g., sparse autoencoders) with circuit-level interventions (e.g., path patching) to uncover representation-mechanism interplay. iii.) Integrating intrinsic interpretability approaches with post-hoc analysis for robust understanding.
For example, coordinated methods could be used for reverse engineering trojaned behaviors (Casper et al., 2023c), where observational techniques identify suspicious activations, interventional methods isolate the relevant circuits, and intrinsic approaches guide the design of more robust architectures.
@@
Another concerning pattern is the tendency to cherry-pick results, relying on a small number of convincing examples or visualizations as the basis for an argument without comprehensive evaluation (Räuker et al., 2023). This amounts to publication bias, showcasing an unrealistic highlight reel of best-case performance. Relatedly, many interpretability techniques are primarily evaluated on small toy models and tasks (Chughtai et al., 2023; Elhage et al., 2022b; Jermyn et al., 2022; Chen et al., 2023b), risking missing critical phenomena that only emerge in more realistic and diverse contexts. This focus on cherry-picked results from toy models is a form of streetlight interpretability (Casper, 2023), examining AI systems under only ideal conditions of maximal interpretability.
@@ 7.2 Technical Limitations
Analyzing Models Embedded in Environments.
Real-world AI systems embedded in rich, interactive environments exhibit two forms of in-context behavior that pose significant interpretability challenges beyond understanding models in isolation. Externally, models may dynamically adapt to and reshape their environments through in-context learning from the interactions and feedback loops with their external environment (Leahy, 2023). Internally, the hydra effect demonstrates in-context reorganization, where models flexibly reorganize their internal representations in a context-dependent manner to maintain capabilities even after ablating key components (McGrath et al., 2023). These two instances of in-context behavior – external adaptation to the environment and internal self-reorganization – undermine interpretability approaches that assume fixed circuits. For models deeply embedded in rich real-world settings, their dynamic coupling with the external world via in-context environmental learning and their internal in-context representational reorganization make strong interpretability guarantees difficult to attain through analysis of the initial model alone.
Adversarial Pressure Against Interpretability.
As models become more capable through increased training and optimization, there is a risk they may learn deceptive behaviors that actively obscure or mislead the interpretability techniques meant to understand them. Models could develop adversarial "mind-reader" components that predict and counteract the specific analysis methods used to interpret their inner workings (Sharkey, 2022; Hubinger, 2022). Optimizing models through techniques like gradient descent could inadvertently make their internal representations less interpretable to external observers (Hubinger, 2019b; Fu et al., 2023b; von Oswald et al., 2023). In extreme cases, a highly advanced AI system singularly focused on preserving its core objectives may directly undermine the fundamental assumptions that enable interpretability methods in the first place.